OpenStudy (anonymous):

A 95% confidence interval for the mean diastolic blood pressure in women is 62 to 110. What is the margin of error?

ganeshie8 (ganeshie8):

mean = (62+100)/2 = 81

ganeshie8 (ganeshie8):

margin of error = 81-62 = ?

OpenStudy (anonymous):

19

ganeshie8 (ganeshie8):

Yep!

OpenStudy (anonymous):

But my prof did it some other way and got a different answer. She got 24 as margin of error

OpenStudy (anonymous):

and 38 for point of estimation

ganeshie8 (ganeshie8):

I'm not sure how your teacher got those numbers

ganeshie8 (ganeshie8):

Hey looks I made a typo I took the upper bound as 100 instead of 110